# Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting

### v1

@@ Abstract
Figure 2: Overview of Evaluation Benchmarks. We evaluate our method across simulation and real-world tasks. Top (Simulation): We construct Personalized-SIMPLER (left/middle) and Personalized-VLABench (right) by repopulating existing environments with user-specific assets. Bottom (Real-world): We conduct physical experiments on a SO-101 arm using 8 diverse object categories, covering both selection and pick-and-place tasks. In all scenarios, the agent must identify a specific target instance among visually similar distractors, requiring precise instance-level grounding beyond generic category recognition.
@@ 1 Introduction
“Robot, bring my cup.” You may ask for your coffee mug while getting ready for work, or have the robot fetch your pet’s favorite toy from the living room floor. As illustrated in Figure Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting, such scenarios where a user’s personal belongings are mixed with visually similar objects will be a frequent necessity in daily life. To truly assist the user, the robot must discern the subtle visual details that distinguish a specific possession from other counterparts. However, general-purpose policies, such as Vision-Language-Action (VLA) models, typically fail to meet this requirement. Since user-specific instances and the personalized instructions used to refer to them (e.g., “my cup”) are rarely covered during training, these models frequently collapse to category-level recognition and mis-ground the intended instance among same-category distractors.
@@
Recent advancements in robot learning have empowered general-purpose policies to execute diverse manipulation tasks specified by natural language commands (rt22023rt2; kim2024openvla; black2025pi05). By training on large-scale robot datasets (oneill2023openx), these models achieve strong generalization to generic instructions (e.g., “pick up the cup”). However, relying solely on language creates a fundamental bottleneck for personalized assistance. Natural language inherently abstracts rich visual details into broad semantic categories. A standard VLA interprets a specific request for “my cup” simply as “a cup,” effectively ignoring instance-level distinctions. This generalization leads the model to discard specific nuances, such as a unique pattern or a chip on the rim, that are critical for identifying the target among distractors. Even with detailed descriptions, text-only disambiguation remains brittle for VLAs because fine-grained instance cues must be grounded indirectly through language, as we demonstrate in Sec. 5. Moreover, per-object fine-tuning would require additional data and optimization for each personal item, which is impractical at deployment.
@@
To bridge this semantic gap, we formulate the challenge of manipulating personal objects. Our goal is to adapt a frozen VLA to recognize and control a user’s unique item unseen during training, utilizing only a handful of reference images. Specifically, given a small set of reference views (e.g., 4–5 images), the agent must discern the specific target from visually similar distractors in a cluttered scene. This formulation targets precise instance-level generalization, extending VLA capabilities beyond broad category-level recognition.
@@
We address this challenge by equipping frozen VLAs with top-down attention in the cognitive sense, namely a mechanism for allocating focus to task-relevant visual evidence at the input level. We propose Visual Attentive Prompting (VAP), a framework that injects instance-awareness into frozen VLAs. VAP bypasses the need for parameter updates by reformulating personalization as a two-stage inference process: (1) Grounding, where the system identifies the user’s object using reference images, and (2) Visual Prompting, where this grounded knowledge is overlaid directly onto the robot’s observation. By explicitly highlighting the target in pixel space, VAP provides a direct visual focus signal to a frozen policy, enabling it to manipulate the user’s object without relying on language-only disambiguation. As a result, a new personal object can be used immediately after it is registered with a few reference images, without any retraining or fine-tuning.
@@
We validate the efficacy of VAP across a comprehensive suite of simulation and real-world scenarios. Since standard VLA benchmarks primarily evaluate category-level generalization, we establish two new benchmarks, Personalized-SIMPLER and Personalized-VLABench, designed to rigorously test instance-level identification among distractors (Figure 2, Top). Crucially, to demonstrate practical applicability, we extend this evaluation to a physical robot setup, tasking the agent with manipulating diverse personal objects in real-world (Figure 2, Bottom). Our extensive experiments show that VAP significantly outperforms both generic policies and other personalization baselines, confirming that visual attentive prompting is a robust strategy for bridging the gap between semantic understanding and instance-level control.
@@
    Personal Object Manipulation: We introduce a personalization task where an agent must manipulate user-specific objects among visually similar distractors using only a few reference images.
@@
    Visual Attentive Prompting (VAP): We propose VAP, a training-free framework that adapts frozen VLAs by translating personal concepts into visual prompts for precise instance-level control, enabling immediate use upon registering new objects.
@@
Figure 3: VAP builds a memory from a few reference images and grounds the target object with frozen detection and segmentation modules. It then highlights that object and rewrites the instruction for a frozen VLA, tracking it over subsequent frames and enabling training-free personalized actions.
@@ 2 Related Work
VLA Models. Recent advancements have established VLA models as a dominant paradigm for general-purpose robotic manipulation (kawaharazuka2025vla-survey; wen2024diffusion). By integrating large language models with visual encoders, VLAs such as RT-2 (rt22023rt2), OpenVLA (kim2024openvla), Octo (octo2024), GR00T N1.5 (gr00t1.5), $\pi_{0}$ (black2024pi0), $\pi_{0.5}$ (black2025pi05) and $\pi_{0.6}^{*}$ (pi06) enable a single policy to execute diverse instructions. These models rely on massive robot datasets (oneill2023openx; droid2024; walke2023bridgedata; rh20t2024) to generalize to novel objects and semantic commands. However, recent benchmarks reveal that VLAs often fail at fine-grained visual grounding and degrade significantly under distribution shifts (fang2025intact; zhang2025vlabench), In personalization, this makes it hard to identify the correct instance among lookalikes and to handle previously unseen personal objects for user-specific references. In this work, we investigate adapting these frozen backbones to personalized scenarios without architectural modifications.
Personalization in Foundation Models. Personalization has emerged as a central theme across foundation models, aiming to adapt large-scale priors to specific user needs. In Large Language Models (LLMs), work has focused on personalized adapters (tan2024personalizedpeft) and long-term memory banks (zhong2023siliconfriend) that align behavior with individual preferences. In the visual domain, text-to-image personalization spans token optimization methods such as Textual Inversion (gal2022textualinversion), weight tuning techniques like DreamBooth (ruiz2023dreambooth) and Custom Diffusion (kumari2023customdiffusion), and more recent encoder-based approaches such as IP-Adapter (ye2023ipadapter) and PhotoMaker (li2024photomaker) for instant customization. Methods like PerSAM (zhang2023persam) extend this capability to segmentation models, enabling SAM to localize user-designated concepts from a single reference image. Similar ideas have been explored in Vision-Language Models (VLMs), where approaches such as MyVLM (alaluf2024myvlm) and Yo’LLaVA (nguyen2024yollavapersonalizedlanguagevision) apply concept learning to personalized dialogue. Collectively, these works establish personalization as a standard capability in generative models, yet its use remains largely confined to static generation or text-based reasoning, leaving a gap in the embodied control.
@@
Personalization for Robotic Manipulation. Extending personalization to embodied agents introduces additional challenges, as user-specific information must be integrated directly into the robot’s decision-making. Recent works primarily adapt instruction modalities or high-level planning. VLAS (zhao2025vlas) incorporates speech recognition to handle user-specific voice commands. TidyBot (wu2023tidybot) and ProVox (grannen2025provox) customize task scheduling to user preferences, while MEMENTO (kwon2025memento) studies how agents exploit episodic memory for user-specific object semantics and behavioral patterns. Relatedly, PIN (barsellotti2024personalized) conditions an agent on reference images (and a short textual description) of a particular instance to navigate to the target. However, these approaches largely overlook the control gap: enabling a general-purpose policy to physically manipulate a specific target instance (e.g., “my cup”) among visually similar distractors using fine-grained visual cues. Our work addresses this gap by using reference images to construct test-time visual prompts that let a frozen, language-conditioned VLA manipulate the correct instance under personalized commands.
Visual Prompting and Grounded Priors. Visual prompting modifies input images (e.g., with bounding boxes, masks, or visual markers) to guide the attention of vision-language models toward specific regions (Shtedritski23; yang2023setofmark). In robotics, this idea has proven effective for injecting high-level priors into control through explicit visual cues. Methods such as RoboGround (huang2025roboground), MOKA (liu2024moka), PIVOT (nasiriany2024pivot), and TraceVLA (zheng2024tracevla) leverage visual prompts to improve control or reasoning, whereas we use reference images to resolve user-specific object references, i.e., to select the correct object among lookalikes (e.g., “my cup”). Concretely, we ground the user’s object from a few reference images and highlight it in pixel space, enabling a frozen VLA to reuse its generic manipulation skills while acting on the correct personal instance without per-object fine-tuning.
@@ 3 Visual Attentive Prompting
We propose a simple and effective method, Visual Attentive Prompting (VAP), to enable VLA models to manipulate personal objects.
@@ 3.1 Problem Formulation
We consider a standard VLA policy $\pi_{\text{VLA}}(a\mid x,\ell)$ that maps an observation $x$ and a language instruction $\ell$ to an action $a$. The observation $x=(I,s)$ consists of RGB images $I=\{I^{(v)}\}_{v=1}^{V}$ from $V$ cameras and the robot’s proprioceptive state $s$. In our setting, the user provides a small reference set $R_{o}=\{I_{o}^{(k)}\}_{k=1}^{K}$ (where $K\approx 5$) capturing a specific instance $o$ that is unseen during training, while its coarse category (e.g., a cup) may be seen. At inference time, the robot confronts a scene containing the target instance alongside visually similar distractors. Our goal is to generate a modified observation $\tilde{x}$ and instruction $\tilde{\ell}$ that enable the frozen $\pi_{\text{VLA}}$ to identify and manipulate the specific target $o$ using only the reference set $R_{o}$, without any additional training.
@@ 3.2 Overview of VAP
VAP acts as an input-side perceptual adapter: it leaves the VLA parameters untouched and only preprocesses the visual and language inputs. Instead of changing the policy, it modifies what the model sees and how the instruction is phrased. In this way, a generic manipulation policy can behave as if it understood personalized commands. Intuitively, VAP first decides which pixels correspond to the user’s object in the scene, then reshapes the visual and language inputs so that this region becomes the natural focus of the policy.
Formally, at timestep $t$ the robot receives its usual observation $x_{t}=(I_{t},s_{t})$ and a language instruction $\ell$, and it is given a small reference set $R_{o}$ that shows the user’s object $o$ from a few viewpoints. VAP uses these references to construct a prompted observation $\tilde{x}_{t}$ and instruction $\tilde{\ell}$ that direct the frozen VLA $\pi_{\text{VLA}}$ toward the correct object. We describe this process with a tracking-aware grounding function $g$ and a prompting function $p$:
@@
To ground the user’s intent in the current scene, VAP treats the reference set $R_{o}$ as a non-parametric visual memory of the personal object. At execution time, it compares this memory to the live images $I_{t}$ and selects the pixels that best match the user’s object, producing a mask $M_{t}$ in each camera view. This mapping turns a personalized reference such as “my cup” into a concrete region in image space.
[…]
@@ 3.3 Grounding and Prompting Details
We next describe how the grounding function $g$ and prompting function $p$ are instantiated in practice. Our design uses a coarse-to-fine perception pipeline that identifies the user’s object across time and multiple camera views.
Grounding. We keep a memory $\mathcal{M}$ of user-specific objects. Offline, the user provides a small set of reference images for each personal item. A frozen visual encoder $f(\cdot)$ embeds each reference view $I_{o}^{(k)}$ of object $o$, producing vectors $Z_{o}=\{f(I_{o}^{(k)})\}_{k=1}^{K}$. At run time, we load $Z_{o}$ for the requested object and use embeddings to distinguish the target from generic instances of the same category.
At the start of an episode, we parse the personalized instruction (e.g., “bring my cup”) and extract the generic category name $c$ (“cup”). We use $c$ as a text query for an open-vocabulary detector and run it on each camera view $I_{t}^{(v)}$. We obtain a set of category-level bounding box proposals $B_{t}^{(v)}=\{b_{i}\}_{i=1}^{N_{v}}$ for view $v$. This stage finds plausible instances of the category, and personalization happens in the following matching step.
Given the proposals $B_{t}^{(v)}$ and the reference embeddings $Z_{o}$, we retrieve the box that best matches the user’s object. For each candidate $b_{i}\in B_{t}^{(v)}$, we crop $I_{t}^{(v)}[b_{i}]$, embed it with $f(\cdot)$ to obtain $e_{i}$, and compute cosine similarities $\cos(e_{i},z_{k})$ for all $z_{k}\in Z_{o}$. To aggregate evidence across reference views, we ensemble per-reference matches to select the target box:
@@
Each reference view $k$ votes for its most similar proposal, and we select the proposal with the largest vote count, filtering distractors that match the target from the view $v$.
@@ 4 Benchmarks
We evaluate VAP on a suite of simulation and real-world benchmarks that instantiate the personalization setting introduced in Figure 2. These benchmarks ask a frozen VLA to act on a user-specific object instance, given a few reference images with same-category distractors. Table 1 summarizes basic statistics.

### v2

@@ Abstract
Figure 2: Overview of Evaluation Benchmarks. We evaluate in simulation (Personalized-SIMPLER, Personalized-VLABench) and on a real SO-101 robot. In each benchmark, one object is replaced by a user-specific instance, same-category distractors are added, and the policy must ground the correct instance from a few reference images.
@@ 1 Introduction
“Hey robot, bring my cup.” You may ask for your coffee mug while getting ready for work, or have the robot fetch your pet’s favorite toy from the living room floor. As illustrated in Figure Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting, such scenarios where a user’s personal belongings are mixed with visually similar objects will be a frequent necessity in daily life. To truly assist the user, the robot must discern the subtle visual details that distinguish a specific possession from other counterparts. However, general-purpose policies, such as Vision-Language-Action (VLA) models, typically fail to meet this requirement. Since user-specific instances and the personalized instructions used to refer to them (e.g., “my cup”) are rarely covered during training, these models frequently collapse to category-level recognition and mis-ground the intended instance among same-category distractors.
@@
Recent advancements in robot learning have empowered general-purpose policies to execute diverse manipulation tasks specified by natural language commands (Zitkovich et al., 2023; Kim et al., 2024; Intelligence et al., 2025). By training on large-scale robot datasets (Open X-Embodiment Collaboration et al., 2023), these models achieve strong generalization to generic instructions (e.g., “pick up the cup”). However, relying solely on language creates a fundamental bottleneck for personalized assistance. Natural language inherently abstracts rich visual details into broad semantic categories. A standard VLA interprets a specific request for “my cup” simply as “a cup,” effectively ignoring instance-level distinctions. This generalization leads the model to discard specific nuances, such as a unique pattern or a chip on the rim, that are critical for identifying the target among distractors. Even with detailed descriptions, text-only disambiguation remains brittle for VLAs because fine-grained instance cues must be grounded indirectly through language, as we demonstrate in Sec. 5. Moreover, per-object fine-tuning would require additional data and optimization for each personal item, which is impractical at deployment.
@@
To bridge this semantic gap, we formulate the challenge of manipulating personal objects. Our goal is to adapt a frozen VLA to recognize and control a user’s unique item unseen during training, utilizing only a handful of reference images. Specifically, given a small set of reference views, the agent must discern the specific target from visually similar distractors in a cluttered scene. This formulation targets instance-level generalization, extending VLA capabilities beyond broad category-level recognition.
@@
We propose Visual Attentive Prompting (VAP), a framework that injects instance-awareness into frozen VLAs. We frame personalization not as model adaptation, but as inference-time input optimization. VAP employs visual interventions to provide an explicit input-side conditioning signal, aligning personal referents with generic cues that frozen VLAs can reliably exploit. To achieve this, VAP reformulates personalization as a training-free two-stage inference process: (1) Grounding, where the system identifies the user’s object using reference images, and (2) Visual Prompting, where this grounded knowledge is overlaid directly onto the robot’s observation. By explicitly highlighting the target in pixel space, VAP provides a direct visual focus signal to a frozen policy, enabling it to manipulate the user’s object without relying on language-only disambiguation. As a result, a new personal object can be used immediately after it is registered with a few reference images, without any retraining or fine-tuning.
@@
We validate the efficacy of VAP across a comprehensive suite of simulation and real-world scenarios. Since standard VLA benchmarks primarily evaluate category-level generalization, we establish two new benchmarks, Personalized-SIMPLER and Personalized-VLABench, designed to rigorously test instance-level identification among distractors (Figure 2, Top). Crucially, to demonstrate practical applicability, we extend this evaluation to a physical robot setup, tasking the agent with manipulating diverse personal objects in real-world (Figure 2, Bottom). Our experiments show that instance-level personalization for frozen VLAs can be achieved by an input-side adapter that couples reference-based grounding with an aligned visual highlight and instruction rewrite. Ablations confirm that neither component alone reliably closes the gap between semantic commands and instance-level control.
@@
    Personal Object Manipulation: We introduce a personalization task for VLAs where the policy must manipulate user-specific objects among visually similar distractors using only a few reference images.
@@
    Visual Attentive Prompting (VAP): We propose VAP, a simple-yet-effective input adapter that couples reference-based instance grounding with a mask-aligned highlight and a matched instruction rewrite, enabling frozen VLAs to act on user-specific instances.
@@
Figure 3: VAP builds a visual memory from a few reference images, grounds the target with frozen detection and segmentation, and prompts a frozen VLA by highlighting the object and rewriting the instruction. A tracker propagates the mask across frames, enabling training-free personalized manipulation.
@@ 2 Related Work
VLA Models. VLA policies have emerged as a practical approach for general-purpose robotic manipulation (Kawaharazuka et al., 2025; Wen et al., 2025; Zitkovich et al., 2023; Kim et al., 2024; Black et al., 2024; Intelligence et al., 2025; Physical Intelligence and others, 2025). Trained on large-scale robot datasets (Open X-Embodiment Collaboration et al., 2023; Khazatsky et al., 2024; Walke et al., 2023; Fang et al., 2023), they can follow open-ended semantic commands and generalize to novel objects. However, recent benchmarks show that VLAs remain brittle under distribution shift and often struggle with fine-grained visual grounding (Fang et al., 2025; Zhang et al., 2025). This limitation is amplified in personalization, where the agent must identify a specific instance among lookalikes and act on previously unseen personal objects. We therefore study how to adapt frozen VLA backbones to personalized scenarios without modifying their parameters.
Personalization in Foundation Models. Personalization seeks to adapt large-scale priors to specific user contexts via tuning or conditioning. In LLMs, this involves lightweight adapters (Tan et al., 2024) or memory banks (Zhong and others, 2024). Similarly, text-to-image customization spans token optimization (Gal et al., 2023), weight tuning (Ruiz et al., 2023; Kumari et al., 2023), and encoder-based projection (Ye et al., 2023; Li et al., 2024), extending even to segmentation (Zhang et al., 2024) and dialogue (Alaluf et al., 2024). However, these methods are predominantly confined to static generation or textual reasoning. They do not address the demands of embodied control, where an agent must not merely generate or describe a personal object, but physically interact with it in a dynamic environment.
@@
Personalization for Robotic Manipulation. Prior personalization efforts have primarily targeted instruction modalities or high-level planning. For instance, VLAS (Zhao et al., 2025) incorporates speech recognition for voice commands, while TidyBot (Wu et al., 2023) and ProVox (Grannen et al., 2025) customize task scheduling. MEMENTO (Kwon et al., 2025) utilizes episodic memory to retrieve object semantics, and PIN (Barsellotti et al., 2024) conditions on reference images for navigation. However, these approaches do not address the control gap: enabling a general policy to physically manipulate a specific instance (e.g., “my cup”) amongst visual distractors. We bridge this gap by constructing test-time visual prompts from reference images, allowing a VLA to reuse generic skills while manipulating the correct personal instance.
Visual Prompting and Grounded Priors. Visual prompting steers vision-language models by overlaying explicit cues, such as bounding boxes or masks, onto input images (Shtedritski et al., 2023; Yang et al., 2023). In robotics, recent works (Huang et al., 2025; Liu et al., 2024a; Nasiriany et al., 2024; Zheng et al., 2025) utilize these markers to inject spatial priors for improved control. In contrast, we leverage visual prompting for instance disambiguation. Instead of enhancing general manipulation capabilities, we use prompts to ground user-provided references, enabling a VLA to identify and act on specific personal objects among visually similar distractors.
@@ 3 Visual Attentive Prompting
We propose Visual Attentive Prompting (VAP), a simple-yet-effective mechanism that steers VLA policies to manipulate user-specified instances without additional training.
@@ 3.1 Problem Formulation
We consider a pre-trained VLA policy $\pi_{\text{VLA}}(a\mid x,\ell)$ mapping observation $x=(I,s)$ and instruction $\ell$ to action $a$. Here, $I=\{I^{(v)}\}_{v=1}^{V}$ denotes multi-view RGB images from $V$ cameras and $s$ represents the proprioceptive state. The user specifies a target instance $o$ via a small reference set $R_{o}=\{I_{o}^{(k)}\}_{k=1}^{K}$ (with $K\approx 5$). Crucially, while the coarse category of $o$ (e.g., a cup) is known, the specific instance is novel and unseen during training. At test time, the robot encounters $o$ amidst visually similar distractors.
[…]
@@ 3.2 Overview of VAP
VAP acts as a training-free inference framework that intervenes strictly in the input space. It first grounds the user’s target instance in the scene using reference images, and then reshapes the visual and language inputs to provide an explicit pixel-level conditioning signal for the target region. This approach enables a generic manipulation model to identify and act on specific personal objects without any weight updates.
Formally, at timestep $t$, the robot receives its observation $x_{t}=(I_{t},s_{t})$, language instruction $\ell$, and a reference set $R_{o}$ depicting the user’s object $o$. VAP uses these references to construct a prompted observation $\tilde{x}_{t}$ and instruction $\tilde{\ell}$ that direct the frozen VLA $\pi_{\text{VLA}}$ toward the correct object. We formulate this process with a tracking-aware grounding function $g$ and a prompting function $p$:
@@
To ground user intent, VAP utilizes $R_{o}$ as a non-parametric visual memory. It matches the memory against the live images $I_{t}$ and selects the pixels that best match the user’s object, producing a mask $M_{t}$ in each camera view. This process effectively maps an abstract reference (e.g., “my cup”) into a concrete region in image space.
Once localized, VAP uses this mask to provide an explicit visual focus cue to the frozen policy. It modifies the observation by overlaying a semi-transparent highlight on the target region while preserving the background context. Simultaneously, it rewrites the instruction into a generic expression aligned with this visual cue (e.g., “pick up my cup” $\rightarrow$ “pick up the red object”). The frozen policy then executes the task using its generic capabilities, yet with attention explicitly guided to the correct instance, enabling personalized control without parameter updates.
@@ 3.3 Grounding and Prompting Details
We implement the grounding function $g$ via a coarse-to-fine perception pipeline: open-vocabulary detection proposes category-level candidates, reference matching identifies the target instance, and temporal tracking propagates the mask over time. In multi-view settings, we execute this pipeline independently for each camera view. We validate critical design choices—including instruction rewriting, voting aggregation, reference count, and prompt styles—via detailed ablations in Appendix E.
Grounding. We keep a memory $\mathcal{M}$ of user-specific objects. Offline, the user provides a small set of reference images for each personal item. In our benchmarks, we construct object-centric reference crops $R_{o}$ by detecting the target in each reference frame (with the same category-level detector used at test time) and cropping around the detected box to minimize background bias. A frozen visual encoder $f(\cdot)$ embeds each reference view $I_{o}^{(k)}$ of object $o$, producing vectors $Z_{o}=\{f(I_{o}^{(k)})\}_{k=1}^{K}$.
At the start of an episode, we parse the personalized instruction (e.g., “bring my cup”) and extract the generic category name $c$ (“cup”). We use $c$ as the text query for an open-vocabulary detector and run it on each camera view $I_{t}^{(v)}$, obtaining category-level bounding box proposals $B_{t}^{(v)}=\{b_{i}\}_{i=1}^{N_{v}}$. This stage isolates plausible category-level candidates, leaving specific instance identification to the subsequent matching step. If no candidates are detected ($N_{v}=0$), we bypass prompting and pass the unmodified view to the VLA. We apply the same fallback if the tracker output mask is unavailable for a view.
Given the proposals $B_{t}^{(v)}$ and the reference embeddings $Z_{o}$, we retrieve the box that best matches the user’s object. For each candidate $b_{i}\in B_{t}^{(v)}$, we crop $I_{t}^{(v)}[b_{i}]$, embed it with $f(\cdot)$ to obtain $e_{i}$, and compute cosine similarities $\cos(e_{i},z_{k})$ for all $z_{k}\in Z_{o}$. To aggregate evidence across reference views, we aggregate per-reference matches to select the target box:
@@
Each reference view $k$ votes for its most similar proposal, and we select the proposal with the largest vote count, favoring proposals supported consistently across references. If multiple proposals tie in vote count, we select among the tied proposals the one with the highest mean cosine similarity to the reference embeddings.
@@ 4 Benchmarks
We evaluate VAP on simulation and real-world tabletop benchmarks that instantiate the personalization setting in Figure 2. Each benchmark requires a frozen VLA to act on a user-specific object instance given a few reference images, in the presence of same-category distractors. Table 1 summarizes benchmark statistics.